if the distribution of a car’s transmission is governed by exponential density function with mean 100000, what will be the probability that car’s transmission will fail during its first 50000 km of operation.


1
Expert's answer
2021-07-07T11:16:25-0400
"f(x; \lambda)=\begin{cases}
 \lambda e^{-\lambda x} &\text{if } x\geq 0 \\
 0 &\text{if } x<0
\end{cases}"

"\lambda=\dfrac{1}{100000}=10^{-5}"



"P(X\leq50000)=1-e^{-10^{-5}(50000)}=1-e^{-0.5}\approx0.39347"

"\approx0.39347"